# Two Causally Related Needles in a Video Haystack

### v1

@@ Abstract
Evaluating the video understanding capabilities of Video-Language Models (VLMs) remains a significant challenge. We propose a long-context video understanding benchmark, Causal2Needles, that assesses two crucial abilities insufficiently evaluated by existing benchmarks: (1) the ability to extract information from two separate locations in a long video and understand them jointly, and (2) the ability to model the world in terms of cause and effect in human behaviors. Specifically, Causal2Needles introduces 2-needle questions, which require extracting information from both the cause and effect human-behavior events in a long video and the associated narration text. To prevent textual bias, these questions comprise two complementary formats: one asking to identify the video clip containing the answer, and one asking for the textual description of an unrelated visual detail from that video clip. Our experiments reveal that models excelling in pre-existing benchmarks struggle with 2-needle visual grounding, and the model performance is negatively correlated with the distance between the two needles. These findings highlight critical limitations in current VLMs. The dataset is available at: https://huggingface.co/datasets/causal2needles/Causal2Needles
@@ 1 Introduction
On many popular benchmarks (Hendrycks et al., 2020; Cobbe et al., 2021; Srivastava et al., 2022; Chen et al., 2021), recent AI systems achieved performance comparable to humans. However, real-world observations suggest that there is still a significant gap between AI and human capabilities (West et al., 2023; Frieder et al., 2023; Amirizaniani et al., 2024). This apparent contradiction suggests that (1) there is much overfitting to popular benchmarks such as GSM8K (Mirzadeh et al., 2024) through mechanisms like data contamination (Singh et al., 2024), and (2) existing benchmarks may not fully reflect differences between machine intelligence and human intelligence (Dziri et al., 2023; Wu et al., 2024). As a result, the development of sophisticated benchmarks that critically evaluate model capabilities has become a high priority for AI research.
@@
Within the context of long video understanding, we investigate two crucial limitations of existing benchmarks for Video-Language Models (VLMs). First, benchmarks evaluating information extraction from a single location do not fully reflect the long video understanding ability of VLMs. A popular type of evaluation for long-context models is called the “needle in a haystack” problem (Kamradt, 2023; Wang et al., 2024b), where the “needle” represents the information to be extracted from the context. Some long video benchmarks feature 1-needle questions, requiring information from a single location in the video. However, research in NLP indicates that models excelling at 1-needle questions often experience significant performance drops when required to extract multiple needles or understand them jointly, revealing the limitations of 1-needle questions for assessing long-context understanding (Li et al., 2023; Vodrahalli et al., 2024; Yang, 2024; Levy et al., 2024). However, multiple-needle questions remain rare in the multimodal setting.
@@
Second, existing benchmarks offer an incomplete evaluation of whether VLMs possess an internal world model. A world model captures the underlying mechanisms governing object dynamics and human behaviors, enabling predictions of future events given an intervention (Forrester, 1971; Ha and Schmidhuber, 2018). However, existing evaluation of world models focus solely on object motion prediction while neglecting human behaviors (Guan et al., 2024a; Motamed et al., 2025; Kang et al., 2024).
@@
To address these limitations, we propose a long-context video understanding benchmark, Causal2Needles, including 902 1-needle questions and 1,494 2-needle questions. It evaluates two key abilities of VLMs: (1) the ability to extract relevant information from two locations in long videos and jointly reason about them, and (2) the ability to model the world in terms of cause and effect of human behaviors.
@@
To evaluate the first ability, we design 2-needle questions using a bridge entity, or the information shared between two causally related events. The identity of the bridge entity is purposely obscured in the 2-needle question, so that the second needle cannot be retrieved before the identity is resolved, which in turn requires the first needle to be retrieved. This mechanism forces the model to jointly understand the two needles in order to answer the question. As a concrete example, Fig. 1 shows a cause event and an effect event that share a bridge entity, “Superman’s death,” ambiguously referred to as “tragedy.” To answer part 1 of the question, the model must first retrieve the effect event, which reveals the tragedy is Superman’s death. Only after that, the model can answer part 2 by retrieving the clip showing the cause event. Suppose the bridge entity were explicitly stated, the question would become “how did Superman die, leading to his memorial service in Metropolis” and be answerable by retrieving the cause event directly. In that case, joint understanding of two needles would be unnecessary and the 2-needle question would degenerate to a 1-needle question.
To assess the second ability, we design our benchmark to test VLMs on identifying the causes of events across diverse, often fictional scenarios. Causal reasoning is fundamental for understanding how actions lead to outcomes, enabling models to go beyond surface patterns to grasp underlying mechanisms (Pearl et al., 2000). Causal2Needles evaluates the understanding of human behaviors from a causal perspective, complementing existing assessments of world models.
@@
Figure 2: The evaluation framework of Causal2Needles. To help models understand the storyline, we also feed the full textual narration into the model. Three types of questions are designed for each pair of causally related events.
A potential drawback of Causal2Needles is that the video clips could be difficult to understand without the narration, but the narration text could introduce textual bias that allows questions to be answered from text alone without video (Ko et al., 2023; Cores et al., 2024; Xiao et al., 2024). As mitigation, we introduce two complementary question formats. The first format, which we call the visual grounding question, requires the model to select the video clip corresponding to the event it needs to retrieve. The second format, which we call the image description question, requires the model to answer multiple-choice questions about the appearance of the retrieved video clip. Fig. 2 shows some examples. Visual grounding questions necessitate the understanding of video clips, but they may underestimate model performance as the format may be out-of-distribution (OOD) to most VLMs. In contrast, image description questions circumvent the OOD issue, but the VLM may benefit from general knowledge of the movie from pretraining data, such as the color of Lois Lane’s outfit, leading to overestimation of model abilities. By evenly splitting 2-needle questions between the two question formats (747 each), Causal2Needles establish a comprehensive benchmark.
Experiments reveal several important findings. First, models that perform well on 1-needle questions exhibit remarkable performance drops on 2-needle questions. Second, the distance between the two needles is negatively correlated with model performance. Taken together, these results suggest that joint understanding of two separate video locations is an important weakness of existing VLMs. We summarize the contributions of this paper as follows:
$\bullet$ We propose the Causal2Needles benchmark that contains 2-needle questions, generated from story videos. We devise a question creation strategy utilizing bridge entities to force the video-language models to jointly understand two video clips arbitrarily located in the context.
@@ 2.1 Long Video Understanding Benchmarks
The ability to retrieve information from a specific location in a long video (often called the “needle in a haystack” task) is widely regarded as a key performance metric for the long-video understanding ability (Kamradt, 2023; Wang et al., 2024b). In Tab. 1, we compare recent long-video benchmarks based on needle-in-haystack tasks side by side to highlight the unique contribution of our Causal2Needles.
We focus exclusively on benchmarks involving multi-needle questions. EgoSchema (Mangalam et al., 2023) fail to assess specific model capabilities due to the lack of type-wise evaluation (e.g., 1-needle vs. 2-needle, reasoning vs. recognition), resulting in limited diagnostic precision. MLVU (Zhou et al., 2024) require models to retrieve visually distinct images that are artificially inserted into videos, which may allow for shortcut features due to the domain gap between the artificial needle and the original content. In contrast to datasets dominated by single-needle questions, MVBench (Li et al., 2024c) and TVBench (Cores et al., 2024) include a large proportion of temporal tracking questions that evaluate a model’s ability to track the state of an object across multiple temporal locations. A similar type of questions is also featured in LongVideoBench (Wu et al., 2025) (e.g., L2-SSS questions) and VideoMME (Fu et al., 2024) (e.g., Counting problems). These temporal tasks typically require only independent understanding of each needle. Moreover, the relationships between their needles often rely on surface-level visual cues, such as matching object colors or identities.
@@
Benchmark Video Length # QA Diagnostic Precision Needle Type Joint Understanding Causal Reasoning
EgoScheme (Mangalam et al., 2023) 180 s 5,000 Natural
MVBench (Li et al., 2024c) 16 s-40 s 4,000 Natural
TVBench (Cores et al., 2024) 16 s-40 s 2,525 Natural
MLVU (Zhou et al., 2024) 180 s-3600 s 3,102 Artificial
VideoMME (Fu et al., 2024) 1018s 2,700 Natural
LVB Wang et al. (2024a) 473 s 6,679 Natural
Causal2Needles (Ours) 438 s 2,396 Natural
Differing from the datasets above, our benchmark requires models to jointly understand both the cause and the effect—the understanding of one relies on the correct interpretation of the other. Furthermore, the relation between the two needles is causality rather than visual similarity. These designs establish a challenging and unique benchmark for long-video understanding.
@@ 2.2 World Model
There is an ongoing debate (Gurnee and Tegmark, 2023; Liu et al., 2025; Guan et al., 2024b; Hu et al., 2024; Kang et al., 2024; Motamed et al., 2025) over whether deep learning networks develop internal world models—whether they learn the underlying principles governing entities, which can be used to to predict future outcomes of possible interventions (Forrester, 1971; Ha and Schmidhuber, 2018).
However, this debate largely overlooks causality between events involving humans, which could be induced by factors like biology (e.g., heavy sweating causes thirst), psychology (e.g., hearing compliments makes people feel good), and social norms (e.g., providing good service leads to tipping). Although such causal questions occasionally appear in video understanding datasets, for example in the form of episodic reasoning in MVBench, it is usually difficult to isolate them and quantitatively measure this factor independently. In contrast, Causal2Needles is dedicated to evaluating reasoning over cause and effect in the context of human behaviors, which allows precise diagnosis of this capability and fills a gap in benchmarks of world models.
@@ 3 Causal2Needles: Dataset Construction
Causal2Needles is built on two video-language datasets, YMS (Dogan et al., 2018) and SyMoN (Sun et al., 2022). With a total of 192 fully annotated movie recap videos, the datasets offer a rich collection of human-behavior events from diverse movie genres. Each event consists of a narration sentence and its corresponding video clip. Details such as the distribution of movie themes and the temporal distance between causal events can be found in Appendix Sec. A.
Our primary focus is to generate challenging 2-needle questions. The generation process is based on a Large Language Model (LLM), ChatGPT-o1, and is decomposed into three steps. First, we extract event-level causal relationships (sentence pairs) from the narration of a video. Then, we rephrase each pair of sentences to explicitly establish a bridge entity between them. Finally, based on the bridge entity, we generate 2-needle questions requiring a joint understanding of both cause and effect events. Each step is guided by carefully designed prompts and illustrative examples (see Appendix Sec. F.1 for details), enabling the LLM to perform the generation incrementally and reliably. Representative examples of the generated questions are shown in Fig. 2. The three steps are elaborated in detail in Sec. 3.1 and 3.2.
@@ 3.1 Causal Relationships Extraction
We employ an LLM to extract causal relationships from narration by combining global and local event graphs, based on an event graph extraction method (Sun et al., 2024). Each event (a narration sentence) is represented as a node, and causal relationships as directed edges. The global graph is extracted from the complete narration of a video. The graph captures long-range causal relationships but may be incomplete, because LLM often overlook sentences in the middle section of a long context (Liu et al., 2024). To address this, we introduce a sliding-window approach that extracts local graphs from 15-sentence segments with a 5-sentence stride. These local graphs capture more detailed causal relationships with a shorter range. We merge the global and local graphs to obtain comprehensive and long-range causal relationships. To avoid superficial causal relationships resulting from temporal adjacency, we retain only those where the cause and effect are separated by at least three events. In Causal2Needles, the distance between cause and effect events ranges from 3 to 21. Appendix Sec. A has more details.
@@ 3.2 Generation of Two-Needle Questions
The generation of 2-needle questions is based on a bridge entity that connects cause and effect events. However, even if two events have a causal relationship, their corresponding narration sentences may not contain an obvious bridge entity because the original narration does not explicitly emphasize causality. In light of this, we prompt a VLM (Gemini-1.5-pro) to rephrase the cause and effect sentences to explicitly establish the bridge entity. For example, the cause sentence, “As it dies, the creature stabs and kills the weakened Superman with one of its bone protrusions,” and the effect sentence, “A memorial is held for Superman in Metropolis,” are rephrased into sentences in Fig. 2. To maintain semantic consistency between the rephrased sentences and their corresponding scenes, we also input the scenes into the VLM as a constraint.
@@
After obtaining causal sentence pairs with a clear bridge entity, we utilize an LLM to generate two parts of a visual grounding 2-needle question. We use the causal relationship in Fig. 2 to demonstrate this process.
First, we prompt the LLM to generate a Part 1 question that requires retrieving the effect event to identify the bridge entity. Specifically, we ask it to extract the bridge entity that establishes the causal relationship, such as “Superman’s death.” Next, we instruct the LLM to rephrase this bridge entity as an ambiguous term, such as “tragedy.” Finally, using this ambiguous term as a reference, we prompt the LLM to create a question that uses the effect sentence as context and the bridge entity as the answer. In this case, the final output is: “What tragedy was commemorated by the memorial service in Metropolis?”
Second, we prompt the LLM to generate a Part 2 question grounded on Part 1, which requires retrieving the cause event. Specifically, based on the ambiguous term “tragedy”, we instruct the LLM to generate a question whose answer appears exclusively in the cause sentence. The output is: “How did such a tragedy occur?”
Finally, combining the two parts with task instruction, we build a visual grounding 2-needle question: “What tragedy was commemorated by the memorial service in Metropolis, and how did such a tragedy occur? Select the scenes that contain the answers to the question.” To answer such questions, the model must jointly understand cause and effect events and ground the answers on video clips.
@@
Since the likely out-of-distribution visual grounding questions may underestimate the performance of VLMs, we also generate image description questions with text-option multiple-choice format. These questions share the same Part 1 as visual grounding questions, but modify Part 2 to ask about visual details of the cause event.
[…]
However, the generated questions may reveal the bridge entity or include excessive object details, allowing the model to locate the cause scene without truly understanding the Part 1 question. For instance, the initially generated question reveals the bridge entity “Superman’s death” and explicitly states the name “Doomsday”, reduce the difficulty of locating the correct scene.
To mitigate this issue, we prompt an LLM to obscure the bridge entity and the named entity being inquired about. We then obtain the final image description 2-needle question, which begins with the Part 1 question: “What tragedy was commemorated by the memorial service in Metropolis? When the tragedy that causes the memorial service occurs, what color were the energy blasts emanating from the creature? A. Yellow B. Blue C. Red D. White”
3.3 Generation of One-Needle Questions
With the extracted causal relationships, we use an LLM to generate simple 1-needle questions that ask for the cause of a given event. Specifically, we illustrate the process using the causal relationship in Fig. 2. First, we prompt the LLM to generate a question using the effect sentence as context, with the answer found exclusively in the cause sentence. The generated question is: “How did Superman die, leading to his memorial service in Metropolis?” Then, to mitigate textual bias from the input narration, we combine this question with the visual grounding instruction: “Select the scene that contains the answer to the question.”

### v3

@@ Abstract
Properly evaluating the ability of Video-Language Models (VLMs) to understand long videos remains a challenge. We propose a long-context video understanding benchmark, Causal2Needles, that assesses two crucial abilities insufficiently addressed by existing benchmarks: (1) extracting information from two separate locations (two needles) in a long video and understanding them jointly, and (2) modeling the world in terms of cause and effect in human behaviors. Causal2Needles evaluates these abilities using noncausal one-needle, causal one-needle, and causal two-needle questions. The most complex question type, causal two-needle questions, require extracting information from both the cause and effect events from a long video and the associated narration text. To prevent textual bias, we introduce two complementary question formats: locating the video clip containing the answer, and verbal description of a visual detail from that video clip. Our experiments reveal that models excelling on existing benchmarks struggle with causal 2-needle questions, and the model performance is negatively correlated with the distance between the two needles. These findings highlight critical limitations in current VLMs. The project page is available at: https://limiaoyu.github.io/Causal2Needles
@@ 1 Introduction
On many popular benchmarks (mmlu; cobbe2021gsm8k; srivastava2022bigBench; chen2021evaluating_code), recent AI systems achieved performance comparable to humans. However, real-world observations suggest that there is still a significant gap between AI and human capabilities (west2023generative; frieder2023mathematical; amirizaniani2024llms; chinchure2025black). This apparent contradiction suggests that (1) there is much overfitting to popular benchmarks such as GSM8K (mirzadeh2024gsm) through mechanisms like data contamination (singh2024evaluation), and (2) existing benchmarks may not fully reflect differences between machine intelligence and human intelligence (dziri2023faith; wu2024reasoning). As a result, the development of sophisticated benchmarks that critically evaluate model capabilities has become a high priority for AI research.
@@
Within the context of long video understanding, we investigate two crucial limitations of existing benchmarks for Video-Language Models (VLMs). First, benchmarks evaluating information extraction from a single location do not fully reflect the long video understanding ability of VLMs. A popular type of evaluation for long-context models adopts the “needle in a haystack” problem formulation (kamradt2023; wang2024needle), where the “needle” represents the information to be extracted from the long context. Some long video benchmarks feature 1-needle questions that require information from a single location in the video. However, research in NLP indicates that model performances on 1-needle questions are often much higher than performances on questions requiring the extraction and understanding multiple needles, revealing the limitations of 1-needle questions for assessing long-context understanding (li2023loogle; vodrahalli2024michelangelo; yang2024retrieval; levy2024same). However, multiple-needle questions remain rare in the multimodal setting.
@@
Second, existing benchmarks offer incomplete evaluation of whether VLMs possess an internal world model, which captures the underlying mechanisms governing object dynamics and human behaviors and enables predictions of future events from an intervention (forrester1971; ha2018world). Existing evaluations of world models (guan2024world; motamed2025generative_physics_worldmodel; kang2024far_physics_worldmodel) focus solely on object motion prediction, neglecting human behaviors and event causality (Sun2023EventCI).
@@
To address these limitations, we propose a long-context video understanding benchmark, Causal2Needles, comprising 2,606 1-needle questions and 1,494 2-needle questions. Out of these, 902 1-needle questions and all 2-needle questions involve causal, world-model reasoning. Causal2Needles evaluates two key abilities of VLMs: (1) extracting relevant information from two locations in long videos and jointly reasoning about them, and (2) modeling the world in terms of causes and effects of human behaviors.
@@
The causal 2-needle questions are constructed from a pair of cause and effect events. Each question should require the VLM to first retrieve the effect event and then the cause event. To formulate the question, we identify a bridge entity, which is an entity or a piece of information shared by the cause event and the effect event. Part one of the question asks the VLM to resolve the bridge entity by retrieving the effect event. Part two of the question requires the retrieval of the cause of the effect event. As an example, Fig. 1 shows a cause event and an effect event that share a bridge entity, “Superman’s death,” ambiguously referred to as “tragedy.” To answer question part one, the model must resolve the content of the tragedy by retrieving the effect event, which reveals the tragedy is Superman’s death. Only after that, the model can answer question part two by retrieving the video clip showing the cause of Superman’s death. If the bridge entity were explicitly stated, the question would become “how did Superman die, leading to his memorial service in Metropolis”, which can be answered by retrieving the cause event directly. In that case, the 2-needle question would degenerate to a 1-needle question. This novel problem formulation allows Causal2Needles to assess the two-needle ability and the causal reasoning ability together.
@@
Figure 2: The evaluation framework of Causal2Needles. To help models understand the storyline, we also feed the full textual narration into the model. Four types of questions are designed for each pair of causally related events.
Causal2Needles is constructed on top of movie summary videos, containing video clips from the movies and narration text. However, this may create a shortcut that VLMs can exploit. The video clips could be difficult to understand without the narration text, but if we feed the narration text as input, the VLM may answer the question from the text directly without using the video. This is the infamous phenomenon of textual bias (ko2023_llm_are_reasoner_for_video; cores2024tvbench; xiao2024can).
As mitigation, we introduce two complementary question formats. The first format, called visual grounding, requires the model to select the video clip corresponding to the event it needs to retrieve. The second format, called image description, requires the model to answer multiple-choice questions about the appearance of the retrieved video clip. Fig. 2 shows some examples. Visual grounding questions necessitate the understanding of video clips, but they may underestimate model performance as the format may be out-of-distribution (OOD) to most VLMs. In contrast, image description questions circumvent the OOD issue, but the VLM may benefit from knowledge of the movie learned from pretraining data, such as the color of Lois Lane’s outfit, leading to overestimation of model abilities. Causal2Needles adopts the visual grounding format for all 1-needle questions and splits the 2-needle questions between the two formats (747 each).
Experiments reveal several important findings. First, causal questions appear substantially more difficult than noncausal questions. Second, models that perform well on 1-needle questions exhibit remarkable performance drops on 2-needle questions. Finally, the distance between the two needles is negatively correlated with model performance. Taken together, these results demonstrate that joint understanding of and causal reasoning over two separate video locations remain important weaknesses of existing VLMs. We summarize the contributions of this paper as follows:
$\bullet$ We propose the Causal2Needles benchmark that contains causal 1-needle and 2-needle questions, generated from story videos. We devise a question creation strategy utilizing bridge entities to force the video-language models to jointly understand two video clips arbitrarily located in the context window.
@@ 2.1 Long Video Understanding Benchmarks
Retrieval of information from a specific location in a long video (usually called “needle in a haystack”) is widely regarded as a key capability in long-video understanding (kamradt2023; wang2024needle). In Tab. 1, we compare recent long-video benchmarks with multiple-needles-in-haystack tasks side by side to highlight the unique contribution of the proposed Causal2Needles dataset. EgoSchema (mangalam2023egoschema) does not measure specific model capabilities (e.g., 1-needle vs. 2-needle, reasoning vs. recognition) as the questions are not categorized by the required capability, resulting in limited diagnostic precision. MLVU (zhou2024mlvu) require models to retrieve visually distinct images that are artificially inserted into videos, which may allow for shortcut features due to the domain gap between the artificial needle and the original content. Multi-needle questions also exist in MVBench (li2024mvbench) and TVBench (cores2024tvbench), which contain questions to evaluate the model ability to track an object across multiple temporal locations. Similar questions are also featured in VideoMME (fu2024videomme) (e.g., counting problems) and LongVideoBench (e.g., L2-SSS questions) (wu2025longvideobench). However, questions in the above datasets typically require only independent understanding of each needle. Moreover, the multiple needles often only related by surface-level visual cues, such as matching visual appearances of objects, rather than semantic relations like causality.
@@
Benchmark Video Length # QA
Diagnostic
Precision
Needle
Type
Joint
Understanding
Causal
Reasoning
EgoScheme (mangalam2023egoschema) 180 s 5,000 Natural
MVBench (li2024mvbench) 16 s-40 s 4,000 Natural
TVBench (cores2024tvbench) 16 s-40 s 2,525 Natural
MLVU (zhou2024mlvu) 180 s-3600 s 3,102 Artificial
VideoMME (fu2024videomme) 1018s 2,700 Natural
LVB wang2024lvbench 473 s 6,679 Natural
Causal2Needles (Ours) 438 s 4,100 Natural
In contrast, the Causal2Needles benchmark requires models to jointly understand both the cause and the effect—the understanding of the cause relies on the correct interpretation of the clue in the effect. Furthermore, the two needles are related by causality rather than visual similarity, necessitating semantic understanding. These designs establish a challenging and unique benchmark for long-video understanding.
@@ 2.2 World Model
There is an ongoing debate (gurnee2023language_worldmodel; liu2024world; guan2024worldmodel_drive; hu2024_math_case_worldmodel; kang2024far_physics_worldmodel; motamed2025generative_physics_worldmodel) over whether deep neural networks learn to develop internal world models—whether they can identify the underlying principles, such as laws of physics, that govern the observational data they learn from. Such world models, if present, can be used to predict future outcomes of possible interventions, such as kicking a ball or blowing a candle (forrester1971; ha2018world; FLASH-2025).
However, this debate largely overlooks causality between events involving humans, which could be induced by factors like biology (e.g., heavy sweating causes thirst), psychology (e.g., hearing compliments makes people feel good), and social norms (e.g., providing good service leads to tipping). Although causal questions sporadically appear in video understanding datasets, such as the episodic reasoning questions in MVBench, it is usually difficult to isolate them and quantitatively measure this factor independently. In contrast, Causal2Needles is dedicated to evaluating reasoning over cause and effect in the context of human behaviors, which allows precise diagnosis of this capability and fills a gap in benchmarks of world models.
@@ 3 Causal2Needles: Dataset Construction
Causal2Needles is built on two video-language datasets, YMS (yms) and SyMoN (symon). With a total of 192 fully annotated movie recap videos, the datasets offer a rich collection of human-behavior events from diverse movie genres. Each event consists of a narration sentence and its corresponding video clip. Details such as the distribution of movie themes and the temporal distance between causal events can be found in Appendix Sec. A.
To facilitate the evaluation of models of diverse capabilities, Causal2Needles consists of both 1-needle and 2-needle questions, as well as causal and noncausal questions, which have different difficulty levels. The generation of both 1-needle and 2-needle questions depend on pairs of cause and effect sentences, extracted in a causal relationship extraction step (Sec. 3.1). We describe the generation of 1-needle questions in Sec. 3.2 and the generation of 2-needle questions in Sec. 3.3.
@@ 3.1 Causal Relationships Extraction
We employ a Large Language Model (LLM) to extract causal relationships from narration by combining global and local event graphs, based on an event graph extraction method (Sun2023EventCI). Each event (a narration sentence) is represented as a node, and causal relationships as directed edges. The global graph is extracted from the complete narration of a video. The graph captures long-range causal relationships but may be incomplete, because LLM often overlook sentences in the middle section of a long context (liu-etal-2024-lost). To address this, we introduce a sliding-window approach that extracts local graphs from 15-sentence segments with a 5-sentence stride. These local graphs capture more detailed causal relationships with a shorter range. We merge the global and local graphs to obtain comprehensive and long-range causal relationships. To avoid superficial causal relationships resulting from temporal adjacency, we retain only those where the cause and effect are separated by at least three events. In Causal2Needles, the distance between cause and effect events ranges from 3 to 21 events. Appendix Sec. A and B contain more details. The prompt used is shown in Appendix Fig. 9.
3.2 Generation of One-Needle Questions
With the extracted causal relationships, we use an LLM to generate simple 1-needle questions. To facilitate reader understanding, we illustrate the process using the cause and effect events shown in Fig. 2. We first prompt the LLM to generate noncausal one-needle questions, which ask for a detail in either the cause or the effect event. For the cause event in Fig. 2, the generated question is: “Who does the creature stab and kill?” We can similarly generate another question for the effect event. The causal one-needle questions differ from the noncausal question by requiring understanding of the causal relation between the two events. We prompt the LLM to generate a question asking for the cause event of a specific effect. The generated question for the example in Fig. 2 is: “What happened to Superman, leading to his memorial service in Metropolis?” Finally, to mitigate potential textual bias from the input narration, we combine these questions with the visual grounding instruction: “Select the scene that contains the answer to the question.” The prompt for 1-needle question generation is shown in Appendix Fig. 10.
@@ 3.3 Generation of Two-Needle Questions
The generation of 2-needle questions is based on a bridge entity that connects cause and effect events. The bridge entity serves two critical purposes: it provides a hint that drives the model to locate the cause event, so we do not provide the cause event directly; it forces the model to resolve the content of the vaguely phrased bridge entity by locating the effect event, before it can locate the cause event.
However, the bridge entity may not be explicitly mentioned in the original cause and effect sentences. Therefore, we prompt a VLM to rephrase the cause and effect sentences to explicitly establish the bridge entity. For example, the cause sentence, “As it dies, the creature stabs and kills the weakened Superman with one of its bone protrusions,” and the effect sentence, “A memorial is held for Superman in Metropolis,” are rephrased into sentences in Fig. 2. To maintain semantic consistency between the rephrased sentences and their corresponding clips, we also input the video clips into the VLM as a constraint.
@@
After obtaining causal sentence pairs with a clear bridge entity, we utilize an LLM to generate two parts of a visual grounding 2-needle question. We use the causal relationship in Fig. 2 to demonstrate this process. First, we prompt the LLM to generate a Part 1 question that requires retrieving the effect event to identify the bridge entity. Specifically, we ask it to extract the bridge entity that establishes the causal relationship, such as “Superman’s death.” Next, we instruct the LLM to rephrase this bridge entity as a vague reference, such as “tragedy.” After that, using this vague reference, we prompt the LLM to create a question that uses the effect sentence as context and the bridge entity as the answer. In this case, the final output is: “What tragedy was commemorated by the memorial service in Metropolis?” Further, we prompt the LLM to generate a Part 2 question, which requires retrieving the cause event. In the example, with the vague reference “tragedy”, we instruct the LLM to generate a question whose answer appears exclusively in the cause sentence. The output is: “How did such a tragedy occur?” It is important to keep the bridge entity vague in this question, so that the question does not give the effect event away. Instead, the model must retrieve the video clip of the effect event and resolve the bridge entity reference. Finally, combining the two question parts with a task instruction, we complete a visual grounding 2-needle question: “What tragedy was commemorated by the memorial service in Metropolis, and how did such a tragedy occur? Select the scenes that contain the answers to the question.” To answer it, the model must jointly understand cause and effect events and ground the answers on video clips. The prompts are in Appendix Fig. 12.
@@
The visual grounding question format may be out-of-distribution for some models. As a result, it may underestimate their performance. To fix this, we also generate questions in a complementary format: multiple-choice image description questions. These questions share the same Part 1 as visual grounding questions, but modify Part 2 to ask about visual details of the cause event.
With the cause video clip, the cause sentence, and the effect sentence as input, we prompt an LLM to generate a question that asks for an attribute of a visible object in the clip. This question follows the template: “When the event that causes $\langle\texttt{Effect Event}\rangle$ occurs, $\langle\texttt{Image Description Question}\rangle$.” For example, in the clip where Metropolis mourns Superman, the LLM output could be: “When the battle that causes Superman’s death and leads to the memorial service occurs, what color were the energy blasts emanating from Doomsday?” We then further prompt the LLM to generate four challenging answer options based on the question.
However, the generated questions may reveal the bridge entity or include excessive object details, allowing the model to locate the cause clip without truly understanding the Part 1 question. For instance, the initially generated question reveals the bridge entity “Superman’s death” and the name “Doomsday”, which reduces the difficulty of locating the video clip. To mitigate this issue, we prompt the LLM to obscure the bridge entity and the named entity being inquired about. We then obtain the final image description 2-needle question, which begins with the Part 1 question: “What tragedy was commemorated by the memorial service in Metropolis? When the tragedy that causes the memorial service occurs, what color were the energy blasts emanating from the creature? A. Yellow B. Blue C. Red D. White” The prompts are shown in Appendix Fig. 14 and Fig. 15.